The number of girls and boys in the class are
proportional to 1.2 and 1.4 . If the number of
students is between 66 and 80, then find the
number of boys in the class?

Answer by checkley77(12844)

1.2:1.4=6:7
6+7=13 FOR EVERY 6 BOYS THERE ARE 7 GIRLS.
THEREFORE THERE HAS TO BE GROUPS OF 13 IN THE CLASS.
66/13=5.07
5*13=65 TOO SMALL.
80/13=6.15
6*13=78 WITHIN LIMITS GIVEN.
6*6=36 BOYS.
7*6=42 GIRLS.
36+42=78